OpenStudy (anonymous):

Please help with Trig Identities I'm lost. I'll mark the answer I think it is but I'm not sure. simplify (tan^2 theta +1)/(tan^2 theta) csc^2 theta -1 tan^2 theta 1 <--

OpenStudy (dumbcow):

\[\frac{\tan^{2} x +1}{\tan^{2} x} = 1 + \cot^{2} x = \csc^{2} x\]

OpenStudy (anonymous):

\[\tan ^{2}x+1=\sec ^{2}x\] so \[\sec ^{2}x/\tan ^{2}x=1/\sin ^{2}x=cosec ^{2}\]

OpenStudy (anonymous):

Glad I checked it. my answer didnt sound right but its what I got when I divided it. Thanks, you really helped me out.